OpenStudy (anonymous):

What is the solution to the equation? t(7 – 2) = 9 • 7 – 9 • 2 A. –2 B. 2 C. 7 D. 9

OpenStudy (rational):

factor 9 in the right side

OpenStudy (rational):

t(7 – 2) = 9 • 7 – 9 • 2 t(7 - 2) = 9(7 - 2)

OpenStudy (rational):

(7 - 2) cancels both sides

OpenStudy (rational):

that gives u t = 9

OpenStudy (anonymous):

thanks

OpenStudy (rational):

np :)